# bibliography.bib

### v1

@@ Abstract
We consider the problem of one-dimensional function approximation using shallow neural networks (NN) with a rectified linear unit (ReLU) activation function and compare their training with traditional methods such as univariate Free Knot Splines (FKS). ReLU NNs and FKS span the same function space, and thus have the same theoretical expressivity. In the case of ReLU NNs, we show that their ill-conditioning degrades rapidly as the width of the network increases. This often leads to significantly poorer approximation in contrast to the FKS representation, which remains well-conditioned as the number of knots increases. We leverage the theory of optimal piecewise linear interpolants to improve the training procedure for a ReLU NN. Using the equidistribution principle, we propose a two-level procedure for training the FKS by first solving the nonlinear problem of finding the optimal knot locations of the interpolating FKS. Determining the optimal knots then acts as a good starting point for training the weights of the FKS. The training of the FKS gives insights into how we can train a ReLU NN effectively to give an equally accurate approximation. More precisely, we combine the training of the ReLU NN with an equidistribution based loss to find the breakpoints of the ReLU functions, combined with preconditioning the ReLU NN approximation (to take an FKS form) to find the scalings of the ReLU functions, leads to a well-conditioned and reliable method of finding an accurate ReLU NN approximation to a target function. We test this method on a series or regular, singular, and rapidly varying target functions and obtain good results realising the expressivity of the network in this case.
@@ 1 Introduction
Function approximation is one of the major applications of neural nets (NNs), either directly or indirectly, for such applications as function generation, system identification, and the solution of physical problems such as differential equations. For an input vector ${\mathbf{x}}$ to a NN, the output of the NN can be thought of as a function evaluation $y({\mathbf{x}})$ where the corresponding function $y$ is defined on a suitable domain and has the same degree of smoothness as the NN activation functions. The NN output $y({\mathbf{x}})$ is trained to approximate some desired target function value $u({\mathbf{x}})$. A NN can be regarded as a function generator for instance in the context of neural operators [Kovachki2023], and can be used as means for solving both ordinary and partial differential equations [karniadakis2021a], [E_DeepRitz].
@@
Approximation theory has long considered the problem of efficiently approximating a univariate target function $u(x),x\in[0,1]$[devore1993constructive, Powell_1981]. Examples include the use of basis splines, such as interpolation methods, or the Galerkin methods used in the Finite Element Method. Such classical function approximations are usually linear in their coefficients, as the weights ${\mathbf{w}}$ of a given set of basis functions are optimised. As such the function approximation lies in a linear function space. Typically the basis functions themselves are piece-wise polynomials, defined over a fixed set of knots ${\mathbf{k}}$ where the form of the local polynomial approximation changes. These methods have provable accuracy and there exist efficient, and well-conditioned, algorithms for calculating the optimal weights that reliably converge to a unique solution, often either solving linear systems of equations (directly or by iteration), or by a simple quadratic minimisation problem. However, such linear methods often lack accuracy, and may require a large number of coefficients to achieve a good level of approximation for a function with complex, or singular behaviour. In one-dimension significantly greater accuracy, with far fewer numbers of coefficients, can often be achieved by using a nonlinear approximation method [deboor] such as a Free Knot Spline (FKS), adaptive methods [HuangRussel]. In contrast to standard spline approximations, in such nonlinear approximations both the weights ${\mathbf{w}}$ and the knot locations ${\mathbf{k}}$ are optimised. Even though such methods are nonlinear in the coefficients, effective a-priori and a-posteriori error estimates and interpolation error estimates have been established in certain cases (such as the FKS), linked to the regularity of the meshes over which the approximations are defined [HuangRussel]. Neural nets are also nonlinear approximators, with many features in common with a FKS or an adaptive method, and we will make extensive use of this correspondence in this paper.
@@
A key advantage of using a FKS over a standard spline approximation, is that for a fixed number $N$ of knots ${\mathbf{k}}$ a higher accuracy can be achieved by exploiting the extra degrees of freedom given by moving the knots [DeVore_1998], which is of importance when approximating functions with rapid variation and/or singularities. A special case are linear interpolating FKS (IFKS) where the weights are evaluations of the target function $u(k_{i})$ at the knot locations, and as such the optimisation reduces to finding knot locations ${\mathbf{k}}$ only. Convergence results in various norms for the best interpolating FKS $\Pi_{1}u$ to a general target function $u$ as $N\to\infty$ are given in [HuangRussel] in the context of adaptive mesh generation. In particular, the approximation error between a general target function $u$ and a linear interpolating FKS $\Pi_{1}u$ with an optimal choice of $N$ knots satisfies $\|\Pi_{1}u-u\|_{2}=\mathcal{O}(N^{-2})$. Once the optimal knots ${\mathbf{k}}$ of the IFKS are known, the problem of finding the optimal weights of the FKS is then a well conditioned (near linear) problem.
When the knots ${\mathbf{k}}$ are not chosen in an optimal way, such as taking uniform knots for instance, the approximation error to the target function $u$ depends on the smoothness of $u$. For a smooth function $u$ the same order of convergence is obtained for an interpolating FKS with regular knots or optimal knots. For a rapidly varying function good convergence is only seen when the know spacing is smaller than the smallest length scale of the target function. For a singular function $u$, a higher asymptotic rate of convergence is obtained as $N\to\infty$ for optimally adapted knots which results in a reduced level of expressivity for the interpolating FKS $\Pi_{1}u$ with uniform knots.
In contrast to a FKS, univariate NNs approximate a general target function $u(x)$ through a combination of linear operations and nonlinear/semilinear activation functions which results in function approximations nonlinear in their coefficients. NNs are thus harder to analyse than linear approximation methods and their training can be challenging due to the non-convexity of the associated loss functions, and focusing on the easier case of shallow NNs as opposed to deep NNs can give valuable insights, for instance in terms their landscape [Mei2018], their mean-field analysis [Shevchenko2022] and their gradient dynamics [Williams2019]. A key advantage of NNs is that they are nonlinear approximators with provable expressivity, see [grohs_kutyniok_2022] and [DeVore_Hanin_Petrova_2021] for instance for reviews on the theoretical expressivity of deep NNs. By the Universal approximation theorem [HORNIK1989] any continuous function on a compact set can be approximated by a certain neural network, with a continuous activation function, to any accuracy. In principle a NN approximation also has the attractive feature that it can be ’self-adaptive’ [E_DeepRitz], so that it may be able to account for singularities and rapid variations in the target function.
A common example for an activation function is the rectified linear unit (ReLU) activation function, defined as ReLU$(x)=\max(x,0)=(x)_{+}$, which is often used in practice due to its efficiency and simplicity. In addition, a key feature of ReLU NNs is that they result in a globally continuous and piecewise linear input-output relation. They are thus formally equivalent to a piecewise linear free knot spline (FKS), with the ReLU NN and FKS describing the same set of functions. For a ReLU NN, the approximation error can be explicitly linked to the depth $L$ and width $W$ of the network, with exponential rates of (error) convergence with increasing depth $L$ [YAROTSKY2017]. However, even with theoretical expressivity guarantees, this does not guarantee that the optimal solution is found by the training algorithm with the standard $L^{2}_{2}$ loss if the initialisation is distant from the optimal approximation. Examples of the training algorithm failing on even quite regular problems will be given in this paper.
@@
The complementary advantages of FKS and ReLU NNs, and their formal equivalence, motivates us to investigate the relationship between shallow ReLU NNs and splines, and particular the link between the knot points of the FKS and the breakpoints (where the derivative changes discontinuously) of the shallow ReLU NN. Using the spline lens, shallow univariate ReLU NNs have been investigated in terms of their initialization, loss surface, Hessian and gradient flow dynamics [Sahs2022]. A ReLU NN representation for continuous piecewise linear functions has been studied in [He2020], with a focus on continuous piecewise linear functions from the finite element method with at least two spatial dimensions. A theoretical study on the expressive power of NNs with a ReLU activation function in comparison to linear spline-type methods is provided in [ECKLE2019232]. The connection between NNs and splines have also been considered in other contexts, for instance when learning activation functions in deep spline NNs [Bohra2020].
@@ 1.1 Contributions
In this paper, we analyse the problem of function approximation from the perspective of training both an FKS and a shallow ReLU NN. We show the latter problem is poorly conditioned as the network width $W$ increases, and as a result the training process can lead to a poor approximation of the target function. We propose a novel two-level approach to the nonlinear problem of training (using Adam) the univariate FKS approximation (which involves finding the weights ${\mathbf{w}}$ and (free) knots ${\mathbf{k}}$). This problem is firstly tackled by constructing an effective loss function based on the equidistribution principle for the knot locations. Secondly we then solve the provably well conditioned problem of finding the weights. We then show that this approach can be modified to train a shallow ReLU NN. Firstly by finding the breakpoints of the ReLU NN and then the scaling factors of the ReLU functions. For the ReLU NN we show that this problem is ill conditioned for large $W$, but can be made well conditioned by a linear preconditioning of the ReLU NN (effectively transforming it into an FKS).
Using a variety of (regular, singular and rapidly varying) target functions, we demonstrate numerically that our two-stage training procedure with the novel loss function both the FKS and the shallow ReLU NN can be trained reliably in one spatial dimension to give an highly accurate approximation. In particular if the width of either network is $N$ then we see ${\cal O}(1/N^{4})$ convergence of the mean square difference between the approximation and the target function This contrasts with the worse performance of applying a standard training procedure to (for example) the ReLU NN, in which we do not see convergence to an accurate solution for any $N$, or see slow convergence for moderate to large values of $N$ due to ill-conditioning.
@@ 1.2 Outline
The paper is structured as follows: in Section 2 we describe deep and shallow ReLU networks, as well as free knot splines (FKS) as nonlinear function approximators. In Section 3, we show the formal equivalence of shallow ReLU NN and the FKS representations and investigate the conditioning of the problem of training the weights and scaling parameters of the respective networks. Section 4 defines the loss functions we consider based on the mean-squared error and an equidistribution-based approach to determine the knots, as well as their training. Numerical results for the one dimensional approximation for a range of target functions are shown in Section 5 where we investigate the impact of the choice of loss function, the quality of the function approximation, the evolution of the knot locations as well as the conditioning and the convergence of the loss training. Finally, in Section 6 we will draw our conclusions based on the numerical results, and suggest areas of further investigation.

### v2

# Equidistribution-based training of Free Knot Splines and ReLU Neural Networks

@@ Abstract
We consider the problem of univariate nonlinear function approximation using shallow neural networks (NN) with a rectified linear unit (ReLU) activation function. We show that the $L_{2}$ based approximation problem is ill-conditioned and the behaviour of optimisation algorithms used in training these networks degrades rapidly as the width of the network increases. This can lead to significantly poorer approximation in practice than expected from the theoretical expressivity of the ReLU architecture and traditional methods such as univariate Free Knot Splines (FKS). Univariate shallow ReLU NNs and FKS span the same function space, and thus have the same theoretical expressivity. However, the FKS representation remains well-conditioned as the number of knots increases. We leverage the theory of optimal piecewise linear interpolants to improve the training procedure for ReLU NNs. Using the equidistribution principle, we propose a two-level procedure for training the FKS by first solving the nonlinear problem of finding the optimal knot locations of the interpolating FKS, and then determine the optimal weights and knots of the FKS by solving a nearly linear, well-conditioned problem. The training of the FKS gives insights into how we can train a ReLU NN effectively, with an equally accurate approximation. We combine the training of the ReLU NN with an equidistribution-based loss to find the breakpoints of the ReLU functions. This is then combined with preconditioning the ReLU NN approximation to find the scalings of the ReLU functions. This fast, well-conditioned and reliable method finds an accurate shallow ReLU NN approximation to a univariate target function. We test this method on a series of regular, singular, and rapidly varying target functions and obtain good results, realising the expressivity of the shallow ReLU network in all cases. We then extend our results to deeper networks.
@@ 1 Introduction
1.1 Overview
Approximation of a target function $u({\mathbf{x}})$, ${\mathbf{x}}\in\Omega$ is one of the major applications of neural nets (NNs). Indeed, a NN (regardless of its architecture) can be regarded as a function generator for instance in the context of neural operators [18], and can be used as means for solving both ordinary and partial differential equations [16], [8]. In general, for an input vector ${\mathbf{x}}$ to a NN, the output of the NN is a nonlinear function evaluation $y({\mathbf{x}})$ where the corresponding function $y$ is defined on a suitable domain and has the same degree of smoothness as the NN activation functions. NNs approximate the target function $u(\mathbf{x})$ through a combination of linear operations and nonlinear/semilinear activation functions which results in function approximations nonlinear in their coefficients. The NN output $y({\mathbf{x}})$ is trained to approximate $u({\mathbf{x}})$. A key feature of NNs is that they are nonlinear approximators with provable expressivity, see [10] and [5] for instance for reviews on the theoretical expressivity of deep NNs. By the Universal approximation theorem [13] any continuous function on a compact set can be approximated by a certain neural network, with a continuous activation function, to any accuracy. In principle a NN approximation also has the attractive feature that it can be ’self-adaptive’ [8], so that it may be able to account for singularities and rapid variations in the target function. A commonly used activation function is the rectified linear unit (ReLU), defined as ReLU$(x)=\max(x,0)=(x)_{+}$. This is often used in practice due to its efficiency and simplicity. ReLU NNs have a globally continuous and piecewise linear input-output relation. They are thus in one-dimension formally equivalent to a piecewise linear free knot spline (FKS), with the ReLU NN and FKS describing the same set of functions [5]. similar results hold in higher dimensions although the mapping one from one to another is rather more subtle. For a ReLU NN, the approximation error can be explicitly linked to the depth $L$ and width $W$ of the network, with exponential rates of (error) convergence with increasing depth $L$ [22].
However, as we will demonstrate, even with theoretical expressivity guarantees, this does not guarantee that the optimal solution is found by the training algorithm with the standard $L^{2}_{2}$ loss. This is because the resulting approximation problem is both badly conditioned and highly non-convex. As a result when training we may either not see convergence, or convergence to a sub-optimal solution, with errors which are many orders of magnitude worse than they should be. This is either because the optimal approximation with high expressivity is not found, or the training process is far too slow and is terminated too early. Similar behaviour is seen in PINNS used to approximate the solution of differential equations, see for example [11]. In this paper we will draw on results from classical adaptive approximation theory to show how the training approach for such a NN can be improved significantly, both by using a more appropriate loss function to overcome the convexity issues (so that we can find the most expressive approximation), and also by preconditioning the NN representation to overcome the ill-conditioning of the training procedure. By doing so we can achieve the theoretical expressivity of the NN in cases where the usual training method fails to deliver a good approximation.
1.1.1 Classical approximation theory
@@
Classical approximation theory has long considered the problem of efficiently approximating a target function $u({\mathbf{x}}),{\mathbf{x}}\in\Omega$ [7, 19]. Examples include the use of combinations of basis splines with localised support, for function interpolation or least squares approximation, such as the Galerkin methods used in the Finite Element Method. Such classical function approximations are usually linear in their coefficients, as the weights ${\mathbf{w}}$ of a given set of basis functions are optimised and the function approximation lies in a linear function space. Typically the basis functions themselves are piece-wise polynomials (splines), defined over a fixed mesh (in d-dimensions), or more simply a set of knots ${\mathbf{k}}$ in one-dimension, where the form of the local polynomial approximation changes. These methods have provable accuracy [6] and there exist efficient, and well-conditioned, algorithms for calculating the optimal weights that reliably converge to a unique solution, often either solving well-conditioned linear systems of equations (directly or by iteration), or by a simple quadratic minimisation problem. However, such linear methods often lack accuracy, and may require a large number of coefficients to achieve a good level of approximation for a function with complex, or singular behaviour. An important example of such an approximation in one dimension, is the piecewise linear interpolant $\Pi_{1}u$ defined over a uniform mesh with $N$ points. If $u\in H^{2}(\Omega)$ we then have, as $N\to\infty$
$N^{2}\|\Pi_{1}u-u\|_{2}\leq C\|u\|_{H^{2}(\Omega)}.$ (1)
If $u$ is not in $H^{2}(\Omega)$ this estimate fails and we see a sub optimal rate of convergence Even if $u\in H^{2}(\Omega)$ when approximating a rapidly varying function good convergence is only seen when the local knot spacing is smaller than the smallest length scale of the target function,
1.1.2 Free Knot Splines (FKS)
@@
Significantly greater accuracy, with far fewer numbers of coefficients, is achieved by using an (adaptive) nonlinear approximation method [4] such as a Free Knot Spline (FKS) in one-dimension, or adaptive mesh methods [15], [21] in higher dimensions. In such approximations both the weights ${\mathbf{w}}$ and the knot locations ${\mathbf{k}}$/mesh are optimised. Even though such methods are nonlinear in the coefficients, effective a-priori and a-posteriori error estimates and interpolation error estimates have been established linked to the regularity of the meshes over which the approximations are defined [15]. A key advantage of using such an adaptive approximation over a standard spline approximation, is that for a fixed number $N$ of knots/mesh-points ${\mathbf{k}}$ a much higher accuracy can be achieved by exploiting the extra degrees of freedom given by moving the points ${\mathbf{k}}$ [6], [15], [21], which is of importance when approximating functions with rapid variation and/or singularities. A special case are linear interpolating FKS (IFKS) and similar interpolating approximations in higher dimensions. In these approximations the weights are evaluations of the target function $u(k_{i})$ at the knot/mesh-point locations, and as such the optimisation reduces to finding knot/mesh-point locations ${\mathbf{k}}$ only. Tight convergence results in various norms have been developed in a general dimension $d$ for the best interpolating piecewise linear approximation $\Pi_{1}u$ to a general target function $u$ as $N\to\infty$. Summaries are given in [15], [6] and later in this paper. In particular, in one-dimension, the optimal approximation error between a general target function $u$ (even one of limited smoothness) and a linear interpolating FKS $\Pi_{1}u$ with an optimal choice of $N$ knots satisfies
[…]
The estimate (2) is in general much better than that in (1) and applies for functions which are not in $H^{2}(\Omega)$. Whilst the IFKS is, in general, far from the best approximation, the optimal knots for the best IFKS are usually close to those for the best FKS. Training a general FKS approximation then becomes a two stage process. The first being the (nonlinear) problem of finding the optimal knot/mesh points and the second the (usually) well conditioned problem of finding the correct weights. We show in this paper that this combination gives excellent approximations with high accuracy.
1.1.3 Combining NN and classical adaptive approximation
@@
The complementary advantages of FKS and ReLU NNs, and their formal equivalence, motivates us to exploit the relationship between shallow ReLU NNs and splines, and in particular both the direct correspondence between the knot points of the FKS and the breakpoints (where the derivative changes discontinuously) of the shallow ReLU NN, and also between the coefficients of the FKS and the ReLU NN. Using the spline lens, shallow univariate ReLU NNs have been investigated in terms of their initialization, loss surface, Hessian and gradient flow dynamics [20]. A ReLU NN representation for continuous piecewise linear functions has been studied in [12], with a focus on continuous piecewise linear functions from the finite element method with at least two spatial dimensions. A theoretical study on the expressive power of NNs with a ReLU activation function in comparison to linear spline-type methods is provided in [9]. The connection between NNs and splines have also been considered in other contexts, for instance when learning activation functions in deep spline NNs [1].
To motivate this comparison we consider the approximation of the target function
$u(x)=\tanh(100(x-0.25)),\quad x\in[0,1].$
This function has a narrow boundary layer at $x=1/4$ which must be resolved to give a good approximation. We consider various methods of approximating this function, in particular (i) ’regular’ spline and ReLU approximations with fixed uniform knots/breakpoints (ii) a shallow ReLU NN and an FKS trained with the ’standard’ $L_{2}$ loss function (iii) a two-level approach with the ReLU NN trained to find the optimal break points and then the optimal weights (without preconditioning) and similarly a FKS trained to first find the optimal knots and then the optimal weights. In Figure 1 we plot (left) the resulting $L_{2}^{2}$ loss of the trained networks as a function of the number of knots/width of the network. We can see from the this figure that the ReLUand regular spline perform poorly, as does the ReLU NN trained on the $L_{2}$ loss function. The FKS trained on the $L_{2}$ loss function performs better but still sub-optimally with poor convergence for large $N$. The FKS trained using the two-level approach performs essentially optimally with rapid convergence, but the same approach for the ReLU NN gives a much worse result if no preconditioning is used. The difference in approximation error between the ReLU NN approximation and the correctly trained FKS is striking with a difference in error of many orders of magnitude. Similarly the difference between the regular spline error and the FKS error is also striking due to the incorrect knot location. This latter problem is well conditioned. To see the role that ill-conditioning of the ReLU network plays we consider using both a FKS with $N=64$ knots and the formally equivalent shallow ReLU NN of width $W=64$ which has $N$ break points. In Figure 1 (right) we compare the training, using the $L_{2}^{2}$ error, of both in the case where the knots/break points are fixed to the optimal values for the IFKS.
Figure 1: (Left) A comparison of the convergence of the approximation of the target function $u(x)=\tanh(100(x-1/4))$ with various architectures. This shows (green) the poor performance of the ReLU NN trained using the regular $L_{2}^{2}$ loss function, the better, but still poor, performance of the regular spline with the same loss function (green dashed). We see the still poor performance of the ReLU NN and the FKS trained using the standard $L_{2}^{2}$ loss function (blue,blue dashed) We also see the far better performance of the FKS trained with a two level approach (red dashed) but the much poorer performance of the ReLU NN using the two-level approach without pre-conditioning (red). A preconditioned ReLU NN architecture trained in the two-level approach shows similar optimal performance to the FKS. (Right) A comparison of the training times and convergence of the weights of the FKS and a shallow ReLU NN for the target function $u(x)=\tanh(100(x-1/4))$ with $N=64$, following the calculation of the knots of the interpolating FKS. The FKS trains very rapidly, whereas the ReLU NN trains slowly due to ill-conditioning.
It is clear from Figure 1 that with this set of knots the FKS trains very much more quickly, and to a much more accurate approximation than the ReLU NN for this target function. This, is due to the poor conditioning of the ReLU NN approximation problem for large $N$ associated with the condition number of the normal equations of the training process. Only by finding the breakpoints of the ReLU network first and then preconditioning the problem of finding the weights can the same accuracy as the FKS be realised.
@@ 1.2 Contributions
We analyse the problem of function approximation from the perspective of training both a ReLU NN and a FKS/adaptive spline, and hence improve the conditioning, training and accuracy of the ReLU NN architecture
Our initial theoretical analysis considers the most simple case of univariate approximation using a shallow ReLU NN and an FKS. Whilst this is removed from the usual practice of using deep networks on higher dimensional problems, the simplicity of this problem allows us to conduct a careful mathematical analysis, which gives much insight into the numerical results that we obtain for more general (deep and higher dimensional) problems.
We show rigorously that in this special (univariate, shallow) case the ReLU NN approximation problem is poorly conditioned as the network width $W$ increases. As a result the training process is very slow, leading either to a poor approximation of the target function, or simply to no convergence. We propose a novel two-level approach to the nonlinear problem of training (using the Adam optimiser) the univariate FKS approximation with weights ${\mathbf{w}}$ and (free) knots ${\mathbf{k}}$). This problem is tackled by constructing an effective loss function based on the equidistribution principle for the knot locations of the interpolating FKS, and minimising this using the Adam optimiser. We then solve the provably well conditioned problem of finding the weights and (close-by) knots of the FKS. We then show that this approach can be used to train the shallow ReLU NN, firstly by finding breakpoints of the ReLU NN (as the knots of the interpolating FKS), and then the scaling factors of the ReLU functions. We show that this latter problem can be made well conditioned for large $W$ by a linear preconditioning of the ReLU NN, transforming it into an FKS, and then training the resulting transformed system. Using a variety of (regular, singular and rapidly varying) target functions, we demonstrate numerically that this two-stage training procedure means that the shallow univariate ReLU NN can be trained reliably to give a highly accurate approximation, achieving the theoretical expressivity of these networks.
We then extend our results to univariate approximations using more general activation functions, and also to deeper networks. The complexity of these systems makes an analytical investigation difficult. However we show, by numerical experiments, that the training of the ReLU NN is still ill-conditioned and (if it converges at all) gives a sub-optimal solution. In contrast methods based on spline approximation can be derived which have high accuracy.
@@ 1.3 Outline
The remainder of this paper is structured as follows. In Section 2 we describe univariate ReLU NNs, as well as free knot splines (FKS) as nonlinear function approximators. In Section 3 we show the formal equivalence of shallow ReLU NN and the FKS representations and the links between the breakpoints/knots and the parameters/weights of these respective architectures. Section 4 defines the loss functions often used based on the mean-squared error and discusses the difficulty of optimising the approximation in using these functions to train the knots/breakpoints and weights together. In Section 5 we focus on the problem of finding the optimal knots for an interpolating FKS and introduce a loss function based on equidistribution as a way of guiding the training. In Section 6 we consider the problem of finding the optimum weights having found the knots (as in Section 5). We show that this problem is well-conditioned for the FKS representation of the solution, but is ill-conditioned for the shallow ReLU NN architecture. In Section 7 we then introduce the idea of two-level training for the knots and then the weights of an FKS, and then show how, after preconditioning, it leads to an effective optimisation approach for the shallow ReLU NN. Numerical results for the one dimensional approximation using both a FKS and a shallow network for a range of target functions are shown in Section 8 where we investigate the impact of the choice of loss function, the quality of the function approximation, the evolution of the knot locations as well as the conditioning and the convergence of the loss training. Extensions of this work to different activation functions, deep networks and to higher dimensions are discussed in Section 9. Finally, in Section 10 we will draw our conclusions based on the results, and suggest areas of further investigation.